Compute the total and annual returns on the described investment.

Five years after buying 50 shares of XYZ stock for ​$70 per​ share, you sell the stock for $5200.

1-The total return is: __% (round to one decimal place as needed)

2- the annual return is:__% (round to one decimal place as needed)
Answer #1

1)

Beginning value = 50 * 70 = 3,500

Total return = [(Ending value - beginning value) / beginning value] * 100

Total return = [(5,200 - 3,500) / 3,500] * 100

Total return = 48.57%

2)

Annual return = (Future value / initial investment)1/n - 1

Annual return = (5200 / 3500)1/5 - 1

Annual return = (1.485714)1/5 - 1

Annual return = 1.0824 - 1

Annual return = 0.0824 or 8.24%